the first team of workers recived 50kg of cement less than second team. every hour of work first team was using 150kg of cement while second team used 200kg, in 3 hours the first team had 1.5 times as much remaining cement as second team. how much cement was delivered to each team
Answer by ankor@dixie-net.com(22740)

the first team of workers received 50kg of cement less than second team.
every hour of work first team was using 150kg of cement while second team used 200kg,
in 3 hours the first team had 1.5 times as much remaining cement as second team.
how much cement was delivered to each team.
:
Let a = amt of cement delivered to the 2nd team
then
(a-50) = amt delivered to the 1st team
:
1st team remains = 1.5 times 2nd team remains
(a-50) - 3(150) = 1.5[a - 3(200)]
a - 50 - 450 = 1.5(a - 600)
a - 500 = 1.5a - 900
-500 + 900 = 1.5a - a
400 = .5a
a = 800 kg delivered to the 2nd team
then
800 - 50 = 750 kg for the 1st team
:
:
Check this by find the actual amt each had remaining
Team 1: 750 - 3(150) = 300 kg
Team 2: 800 - 3(200) = 200 kg
300 is 1.5*200